OpenStudy (anonymous):

How can I graph a hyperbola with foci at (0, ±10) and vertices at (0, ±8) using graphing technology.

OpenStudy (jdoe0001):

ok, let's take a peek

OpenStudy (jdoe0001):

vertices are at (0, 8) (0, -8) notice "x" doesn't change, "y" does so is a vertical traverse axis so is going up and down

OpenStudy (jdoe0001):

the distance from the vertices to the (h,k) center, is "a" so, one vertex plus the other, half that, is where the center is in this case is moving up 8, down 8 so the center is clearly (0, 0)

OpenStudy (anonymous):

It's a parabola that opens up and down right? :D

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

parabolas only open either down or up, not both :)

OpenStudy (anonymous):

Alright, so the center would be 0,0? What equation would I use to graph that? Like how do I set up the equation

OpenStudy (jdoe0001):

well, I assume you know what the equation of a hyperbola is?

OpenStudy (anonymous):

It'd be two parabolas then? :D

OpenStudy (jdoe0001):

well, yes, a hyperbola is pretty much 2 parabolas, mirroring each other

OpenStudy (jdoe0001):

you do know what a hyperbola is?

OpenStudy (jdoe0001):

as in, the exercise wouldn't be here if you hadn't covered that already

OpenStudy (jdoe0001):

and I also assume you know a hyperbola equation

OpenStudy (anonymous):

Yeah, that's what we're studying right now. I know how to graph them when I have the equation just I'm not sure how I get the equation when I only have the vertices.

OpenStudy (jdoe0001):

ok, so your (h, k) center is a (0, 0) from there it moves up 8 down 8 to the vertices the distance from the center to either vertex is "a"

OpenStudy (jdoe0001):

so a = 8

OpenStudy (jdoe0001):

now, your foci are at (0, 10) (0, -10) so is going up 10 and down 10 the distance from the center to either focus is "c" so c = 10 now \(c^2 = a^2 + b^2 \) as in pythagorean theorem so \(b^2 = c^2 - a^2\) solve that for \(b^2\) a = 8, so \(a^2 = 64\)

OpenStudy (jdoe0001):

and then put the values in the hyperbola equation :)

OpenStudy (anonymous):

so, I solve \[10^2 = 8^2 + x^2\] Which eould be 2^2?

OpenStudy (amistre64):

it might help to keep your a and b parts under the x and y parts b is conventionally an up down measure, and a is left to right measure. using slope as y/x then b/a conforms nicely to your under xy parts

OpenStudy (anonymous):

Now I'm just confused. :/

OpenStudy (jdoe0001):

amistre64 is correct, which one goes under which one? well "a" always goes to the left-hand-side fraction and "b" to the other :) and since it's a vertical traverse axis hyperbola, up and down the fraction with the POSITIVE sign will the one with the "y"

OpenStudy (jdoe0001):

well, solve for \(x^2\)

OpenStudy (jdoe0001):

so, what does \(10^2 = 8^2 + b^2\) give you for \(b^2\)

OpenStudy (amistre64):

there are 2 parts to an ellips or a hyperbola\[\frac{x^2}{a^2}~and~\frac{y^2}{b^2}~=~1\] since your verts are at (0,\(\pm\)y): (0,\(\pm\)8) \[\frac{0^2}{a^2}~+~\frac{8^2}{8^2}~=~1\] \[-\frac{0^2}{a^2}~+~\frac{8^2}{8^2}~=~1\]rewriting for asthetics \[\frac{y^2}{8^2}-\frac {x^2}{a^2}=~1\]solving a has already been addressed

OpenStudy (anonymous):

a^2 would be 36 so a = 6?

OpenStudy (jdoe0001):

well, a = 8, \(8^2 \ne 36\)

OpenStudy (jdoe0001):

so, what did you get for \(b^2\)?

OpenStudy (anonymous):

oh, is \[b^2 = 36\]

OpenStudy (jdoe0001):

ohh.... I guess you meant b = 6 yes

OpenStudy (jdoe0001):

right, so, keep in mind that is going vertical, so, in a hyperbola you have 2 fractions, one is negative, one is positive the hyperbola goes on the direction of the POSITIVE variable if it's going upwards, that means "y" fraction is POSITIVE

OpenStudy (amistre64):

yes, the "missing" value is 6.

OpenStudy (jdoe0001):

so, how does your hyperbola look like? :)

OpenStudy (jdoe0001):

I guess I was looking for the equation, but yes, that looks right :)